# The Vector Grounding Problem

### v2

@@ 4.1 Referential Grounding
The first notion of grounding, which we call referential grounding, closely aligns with the philosophical notion of reference. [^11] Partee (1981) uses the phrase “language-to-world grounding” to designate a closely related notion, which she contrasts with “language-to-language grounding” (see subsection 4.3 below for discussion of the latter). While Partee’s notion of language-to-world grounding focuses on what ties down the intensions of linguistic expressions with respect to their extra-linguistic content, our notion of referential grounding is broader, including the relation between internal representations and their worldly referents. It is the type of grounding that connects representations to things in the world. Reference can be loosely described as a relation that enables representations to ‘hook onto’ worldly entities or properties. [^12] Much philosophical work has been dedicated to examining the nature and varieties of reference. For our limited purposes, however, the intuitive notion captured by the ‘hooking’ metaphor will suffice. For example, the word ‘Paris’ refers to the city of Paris, and arguably, so does a map of Paris. Both representations allow us to receive and convey information about the actual city of Paris. Similarly, many of our thoughts seem to hook onto specific entities and properties in the world, allowing us to entertain thoughts that refer to Paris, such as beliefs about it being the capital of France or desires to visit the city.
@@ 4.2 Sensorimotor Grounding
Another type of grounding that is often appealed to, especially in cognitive psychology, is what we call sensorimotor grounding. For an internal representation to be grounded in this sense is for it to be fundamentally linked to sensorimotor representations available within a cognitive system. For instance, the lexical representation ‘football’ is sensorimotorically grounded if its representation in a cognitive system can be reduced to, or is intrinsically connected to, sensorimotor representations of kicking movements, of the visual perception of a football pitch, of the auditory perception of the sound of a ball being kicked, etc. This kind of grounding plays a prominent role in theories in cognitive psychology that embrace a moderate form of embodied cognition, such as those positing the existence of “perceptual symbol systems” in human cognition (Barsalou 1999, Pulvermüller 1999).
Interestingly, sensorimotor grounding is also the solution proposed by Harnad (1990) to his Symbol Grounding Problem. He argued that the symbols in classical AI systems needed to be grounded in two types of representation: (a) iconic representations, which depict transformations of the projections of object ‘shapes’ on a system’s sensory surfaces; and (b) categorical representations, which represent invariant sensory features that facilitate the identification and discrimination of objects in different categories. Harnad hypothesised that connectionist approaches to AI would be better equipped to provide the necessary iconic and categorical representations, which could then complement symbolic AI systems and result in hybrid symbolic-connectionist systems.
We will come back to Harnad’s suggestion in subsection 4.6, arguing that it fails to solve the Symbol Grounding Problem, and does not help with the Vector Grounding Problem either.
@@ 4.3 Relational Grounding
A third notion of grounding, that we may call relational grounding, is occasionally used to refer specifically to intra-linguistic relationships, in accordance with the idea that a word’s meaning in a language is partly determined by its relations to other words. [^13] The label is somewhat under-descriptive, since all forms of grounding are relational, though in different ways. What we wish to underline, however, is that on this sense of grounding the meaning of representations depends on the relations they stand in to other representations in the same representational system. Partee (1981) introduces the notion of “language-to-language grounding”, referring to the way in which a language learner acquires knowledge about the meanings of words by learning semantic connections between words in the language. According to Partee, one key way in which this happens is through accepting certain sentences as true, particularly those that are regarded as impervious to empirical challenge within the linguistic community. For example, learning that “bachelors are unmarried males” helps ground the meaning of ‘bachelor’ through its relation to ‘unmarried’ and ‘male’.
Relational grounding is closely related to inferential role semantics (Block 1986, Brandom 1994). Inferential role semantics identifies the meaning of a linguistic expression with its inferential relations to other expressions. For example, the word ‘and’ plays a logical role in reasoning that determines when compound statements are true or false based on the truth or falsity of their components. Likewise, words like ‘larger’ or ‘slower’ have inferential roles in comparisons, allowing us to determine comparisons between items or the relative properties of a single item. Knowledge of such relations between words underlies our ability to draw material inferences in language and thought.
Finally, there is a clear connection between Partee’s “language-to-language grounding” and distributional semantics. Distributional semantics seeks to analyse word meanings in terms of word-to-word relationships. However, instead of drawing on linguistic analyses of word meaning, it relies on analysing co-occurrence statistics in text corpora. Consequently, distributional semantic models capture relationships between words grounded in patterns of use, rather than linguistic theory. The high dimensionality of vector spaces in LLMs makes it possible to represent complex, multifaceted similarities between words – some of which plausibly reflect conceptual structure and/or inferential roles (Piantadosi & Hill 2022).
@@ 4.4 Communicative Grounding
The notion of grounding is also used in relation to another aspect of language, which has to do with communication (Clark & Brennan 1991, Traum 1994). In this context, the ‘grounding problem’ refers to the problem of establishing common ground between speakers in conversation. From this perspective, linguistic communication can be seen as a form of coordinated action that involves collaborating to reach a common understanding of what is said. The process of communicative grounding may involve explicit clarification strategies, such as prompting an interlocutor to repeat or reformulate a garbled or ambiguous statement. It also involves tracking communicative intentions during the exchange, to infer what is communicated from what is said. Such inferences can be made partly on the basis of pragmatic presuppositions, namely, what speakers take for granted or assume to be true when they use certain sentences in the context of the conversation (Stalnaker 2002).
The communicative notion of grounding has also been discussed by natural language processing (NLP) researchers (Brennan 1998, Di Maro 2021, Chandu et al. 2021). For example, Chandu et al. (2021) distinguish between static grounding, in which common ground in basic communication between a human user and a virtual agent is automatically established by querying a database (e.g., asking for a weather report); and dynamic grounding, in which common ground is more gradually established in back-and-forth communication between a human and a dialogue agent. Bender & Koller (2020) define meaning to be a relation between linguistic expressions and communicative intent, and characterise this relation as a form of grounding.
While communicative and referential grounding are distinct notions, they are related in important ways. Referential grounding anchors language to the world, but the same linguistic expressions may be referentially grounded in different ways for different speakers, due to differences in perception, knowledge, and conceptualisation. This means that calibrating reference in conversation requires communicative grounding – language users must work together to negotiate a common way of connecting language to the world. Without such collaboration, there is no guarantee that speakers really mean the same thing, even if they are using the same words.
@@ 4.5 Epistemic Grounding
Finally, the notion of grounding is occasionally used by NLP researchers to refer more minimally to the relationship between linguistic expressions and information stored in knowledge bases (Poon 2013, Tsai & Roth 2016, Chandu et al. 2021, Thoppilan et al. 2022). For example, an expert system may contain a knowledge base of facts and rules belonging to a particular domain. When the system receives user input, it can query this knowledge base to determine an appropriate response. A classic example is MYCIN, an expert system for diagnosing infectious blood diseases (Shortliffe 1977). When given patient symptoms, MYCIN would search its knowledge base of medical rules and facts to suggest possible diagnoses and recommend treatments. In this system, language is grounded in the structured knowledge contained in the expert system’s knowledge base. We call this notion epistemic grounding.
There are two ways in which current LLMs may involve epistemic grounding. Firstly, they can be hooked up to external databases used to retrieve information (Borgeaud et al. 2022, Thoppilan et al. 2022). Secondly, it has been suggested that even LLMs that are not augmented with retrieval capabilities might be functionally similar to knowledge bases themselves (Petroni et al. 2019). Indeed, they may encode a significant amount of world knowledge in their weights after pre-training, as illustrated by the remarkable performance of state-of-the-art LLMs on a wide range of factual question-answering benchmarks (OpenAI 2023, Bubeck et al. 2023).
@@ 4.6 What Grounding for LLMs?
We have identified five distinct notions of grounding that play prominent roles in discussions about meaning grounding in biological and artificial systems. Which of these notions is necessary for LLMs to possess representations and produce outputs endowed with intrinsic meaning? (Recall that intrinsic meaning is meaning that is independent from external interpretation.) To identify the most suitable notion of grounding for this task, we will employ an elimination strategy, examining each notion for its adequacy, beginning with Harnad’s1990 proposal of appealing to sensorimotor grounding as a solution to the Symbol Grounding Problem.
Sensorimotor grounding connects conceptual or linguistic representations to sensorimotor representations within the same cognitive system. This sort of grounding can be useful to shed light on some features of the former representations, how they work, and what roles they can play in the economy of a cognitive system. However, sensorimotor grounding falls short in addressing the merry-go-round problem, as it merely grounds one type of representation in another: it does not tackle the question of how sensorimotor representations establish a connection with the world. One might argue that sensorimotor representations are grounded in the world by resembling, to some extent, the entities they represent. However, the success of such an approach relies on an additional factor beyond sensorimotor grounding itself: it is the relevant resemblance relations that are supposed to ground those sensorimotor representations in the world. As sensorimotor grounding does not provide a connection with the world but only with other representations, it does not address the Symbol Grounding Problem. Similarly, this form of grounding does not resolve the Vector Grounding Problem, as grounding vectors in other vectors still leaves open the question of how any of those vectors come to be connected to the world.
@@
The same line of reasoning applies to relational, communicative, and epistemic grounding. Relational grounding anchors linguistic representations in their relationships to other linguistic representations, suggesting that a representation’s meaning is determined by its connections to the meanings of other representations within a network of inferential, conceptual, or statistical relations. Although this kind of grounding clearly does not escape the representational carousel, inferential-role semantics may suffice for connecting at least some words to what they are about, e.g., logical connectives, as their meanings are arguably exhausted by the inferences they are involved in (Block 1986, Fodor 1990). Logical connectives are however a rather special case, as they are not supposed to represent entities in the world, but rather abstract relations between any possible entities. [^14] Therefore, if an LLM can learn the structure of relations that constitute the meaning of logical connectives and linguistic conjunctions from its training data, then its representations of these terms might be as intrinsically meaningful as they can be. Inferential-role semantics can thereby account, at best, for only a very small and special subset of the representations and outputs that biological and artificial systems possess and produce.
Communicative grounding seeks to anchor representations in the shared content of linguistic interactions between agents. However, this shared content relies on the representations within the agents’ cognitive systems and the meanings of any public representations used in the exchange. As with sensorimotor grounding, one type of representation is grounded in another, perpetuating the representational carousel without establishing a connection to the world beyond the representations.
Epistemic grounding, on the other hand, aims to ground linguistic representations in knowledge bases. Yet, this approach encounters a familiar obstacle: the representations within the knowledge bases themselves call for an explanation of how they acquire intrinsic meaning. Consequently, epistemic grounding alone fails to provide a comprehensive solution to the Grounding Problem, as it does not address the question of how representations come to possess intrinsic meaning to start with.
@@
In sum, the form of grounding that underlies the other four types and is most relevant to addressing the Grounding Problems is what we have called referential grounding. Referential grounding anchors a representation – be it a symbol, a map, or a vector – in the world itself, rather than in other representations. It stands as the sole form of grounding that escapes the representational merry-go-round and connects a representation to whatever worldly entity or property it is about. [^15] Or almost: see the caveat about inferential-role semantics and logical connectives above. Consequently, to determine whether Generative Pre-Trained Models such as LLMs manipulate and produce representations and outputs with intrinsic meaning, we must assess whether they can achieve referential grounding. To do so, we first need to examine the requirements for a representation to be referentially grounded.
@@ 5 Theories of Representational Content: Causes and History
Some internal representations are basic, in the sense that their meaning, or content, does not derive from other states already endowed with meaning. They can be seen as the building blocks of more complex representations, which draw some of their meaning from those basic representations. For instance, ‘edge detectors’ in primary visual cortex (V1) are good candidates for being basic internal representations, whereas ‘shape detectors’ in downstream visual areas are not, since their contents derive partly from the contents of the basic edge detectors. Similarly, words and sentences are not basic in this sense, since their meaning derives from other, more basic representational states, such as beliefs, desires and intentions – themselves complex representations.
The distinction between basic and complex representations cuts across the distinction between intrinsic and extrinsic meaning. As a reminder, states with intrinsic meaning are those that carry meaning independently of an interpreter with mental states, such as a human, attributing meaning to them. Both the representations in primary visual cortex and the words and sentences humans utter are cases of representation with intrinsic meaning, though the former are basic while the latter are complex. By contrast, words randomly formed by the movement of ants in the sand, outputs from a random sentence generator – and, according to Bender & Koller (2020), the outputs of LLMs – only have extrinsic meaning: they are merely shapes that require human interpretation to convey any meaning at all.
We cannot here do justice to the rich and nuanced interdisciplinary research that has been carried out for the past several decades on the notion of cognitive representation. [^16] See Neander (2017), Millikan (2017), Shea (2018) for recent book-length defences of the main contenders today. What matters for our limited purposes is that, despite some disagreements on the finer details, philosophers of cognitive science have largely reached a consensus on the key factors that enable internal states in a system to become representations, both basic and complex (Shea 2018, Coelho Mollo 2022).
@@
Two natural relations are widely considered to be central to endowing basic internal states with meaning: causal-informational relations, and historical relations. In general terms, causal-informational relations occur when one state of the world conveys information about another due to a causally-generated correlation between the two states. A classic example is the relationship between smoke and fire: smoke is correlated with fire because it is caused by it, so the presence of smoke carries information about the presence of fire. Neurons in brain area V1 convey information about edges in the perceived visual scene, as their activation correlates (through a relatively complex causal chain) with the presence of edges in the world in front of the perceiving subject.
@@
Causal-informational relations capture the main rationale for why systems possess basic internal representations: they allow systems to be sensitive to specific features of their environments – those for which internal states stand in causal-informational relations to. However, the mere existence of a causal-informational relation is insufficient to provide an account of basic representational content. An additional requirement often mentioned is that the causal-informational relation be exploited, that is, used by the system due to its information-carrying nature (Shea 2018). In other words, a system must use the presence of smoke to obtain information about the presence of fire, form a belief about there being a fire, guide fire-related behaviour, and so on, for smoke to be a viable candidate for representing fire for that system.
@@
Similarly, the information about the world that edge detectors in V1 carry into the mammalian brain must be used downstream for further processing – such as detecting food sources – so as to guide appropriate behaviour. To be a representation necessarily involves playing a certain role: that of conveying information about the world and being used by the system in virtue of that. This further requirement complements the explanation for why internal representations are useful to systems: they are used by the system as sources of information about the world. It also helps to avoid the risk of trivialising the notion. Causal-informational relations are abundant in nature: every effect carries information about its cause. If every such relation should suffice for being a representation, representations would be so extremely trivial that their special explanatory power for cognition would be lost.
@@
Exploited causal-informational relations alone, however, are still not enough to base an adequate theory of representational content, for another crucial feature of internal representations is missing: their capacity for truth or falsehood, correctness or incorrectness. To be a representation inherently involves the possibility of misrepresenting – the possibility of getting things wrong and failing to fit to the world accurately. V1 neurons may become active when no edge is present, as in the famous Kanizsa’s triangle illusion, thus conveying information to the system that, if not corrected by other subsystems, can lead to inappropriate behaviour (such as trying to touch the non-existent edge). Representations involve a kind of descriptive normativity (Neander 2017). That is to say, representations can represent successfully or unsuccessfully, better or worse. This descriptive normativity extends to external representations, such as sentences, and meaning more generally. They are kinds of things that can be true or false, accurate or inaccurate, appropriate or inappropriate. The rationale for systems to possess representations involves that they be accurate enough, enough of the time. Otherwise, representations would lead systems astray and hinder appropriate behaviour, rather than the other way around. Generally, representations have the function to represent correctly and accurately (to a degree of required precision). [^17] In some special cases, systematic misrepresentation can be a sensible price to pay. If the stakes are high enough, it makes sense for representations to err on the side caution: if you are trying to detect predators, it makes sense to produce many false-positive (misrepresenting predators as present when there are none), rather than even one, potentially fatal false-negative.
@@
As the notion of representation is central for cognitive (and computer) science, the notion of function is central to biology. This has led philosophers of biology to develop scientifically-grounded, objective theories of function Wright (1973), Millikan (1989), Neander (1991), Godfrey-Smith (1994), Garson (2019). If functions exist out there in the world, they must be a matter of natural features and relations, rather than human interpretation. The dominant view of functions is the selected-effects theory Millikan (1989), Neander (1991), Garson (2019). On this view, functions are ensembles of the historical causes that explain the persistence of a certain trait or feature in a system. For instance, hearts today have the function to pump blood because past hearts, by pumping blood, causally contributed to the survival and reproduction of the systems having them. Pumping blood is what explains, causal-historically, why hearts still exist – and it is thereby their function. The relevant causes can be at the level of the species, typically through natural selection; or at the level of the individual, through selection-like learning processes during development.
@@
The most influential views on representation incorporate the selected-effects theory to account for internal states having the function to carry information about the world and to guide appropriate behaviour Neander (2017), Millikan (2017), Shea (2018), Coelho Mollo (2022). The guiding idea is that, through selection and/or learning, states and processes in cognitive systems come to acquire functions they are supposed to perform. For example, neurons in V1 have the function of detecting edges in the visual scene because they were selected for this purpose during evolution via natural selection; and such selection occurred due to the causal-informational relations they maintain with actual edges in the world. When these neurons successfully detect edges, they fulfil their function, representing correctly; when they fail to fire in the presence of an edge or fire in the absence of one, they malfunction, misrepresenting. In brief, internal representations are states that carry information about the world, have the function to do so, and are appropriately used by the system in virtue of the information they carry.

### v3

@@ 4.1 Referential Grounding
The first notion of grounding, which we call referential grounding, closely aligns with the notion of reference. It is the type of grounding that connects representations to things in the world. Reference can be loosely described as a relation that enables representations to ‘hook onto’ worldly entities or properties. [^11] Much philosophical work has been dedicated to examining the nature and varieties of reference. For our limited purposes, however, the intuitive notion captured by the ‘hooking’ metaphor will suffice. For example, the word ‘Paris’ refers to the city of Paris, and arguably, so does a map of Paris. Both representations allow us to receive and convey information about the actual city of Paris. Similarly, many of our thoughts seem to hook onto specific entities and properties in the world, allowing us to entertain thoughts that refer to Paris, such as beliefs about it being the capital of France or desires to visit the city.
@@ 4.2 Sensorimotor Grounding
Another type of grounding that is often invoked, particularly in cognitive psychology, is what we might call sensorimotor grounding. More precisely, we can divide sensorimotor grounding into two related but different views that are often conflated.  [^12] We thank an anonymous reviewer for pressing this point.
According to what we may call the ‘sensorimotor representation’ view of grounding, for an internal representation to be sensorimotorically grounded is for it to be fundamentally linked to sensorimotor representations available within a cognitive system. For instance, the lexical representation ‘football’ is sensorimotorically grounded if its representation in a cognitive system can be reduced to, or is intrinsically connected to, sensorimotor representations of kicking movements, of the visual perception of a football pitch, of the auditory perception of the sound of a ball being kicked, etc. This kind of grounding plays a prominent role in theories in cognitive psychology that embrace a moderate form of embodied cognition, such as those positing the existence of “perceptual symbol systems” in human cognition (Barsalou1999, Pulvermueller1999). This kind of sensorimotor grounding is the approach proposed by Harnad1990 to tackle the Symbol Grounding Problem.
Another possible way to understand sensorimotor grounding is in terms of what we may call the ‘sensorimotor contact’ view of grounding. Its guiding idea is that an organism’s sensory and motor systems interact with the world in a distinctive, unmediated way. Specifically, sensory and motor transducers—such as nerve endings in sensory surfaces—make direct contact with the world, thus providing a basic form of grounding that is inherited, through appropriate causal chains, by sensorimotor systems and organisms as a whole.
@@ 4.3 Relational Grounding
A third notion of grounding, that we may call relational grounding, is occasionally used to refer specifically to intra-linguistic relationships, in accordance with the idea that a word’s meaning in a language is partly determined by its relations to other words. [^13] The label is somewhat under-descriptive, since all forms of grounding are relational, though in different ways. What we wish to underline, however, is that on this sense of grounding the meaning of representations depends on the relations they stand in to other representations in the same representational system. This “language-to-language grounding” (parteeMontagueGrammarMental1981) captures the way in which a language learner can acquire knowledge about the meanings of words by learning semantic connections between words in the language.
Relational grounding is closely related to inferential role semantics (blockAdvertisementSemanticsPsychology1986, brandomMakingItExplicit1994), which identifies the meaning of a linguistic expression with its inferential relations to other expressions. There is also a clear connection between relational grounding and distributional semantics. Distributional semantics seeks to analyse word meanings in terms of word-to-word relationships. However, instead of drawing on linguistic analyses of word meaning, it relies on analysing co-occurrence statistics in text corpora. Consequently, distributional semantic models capture relationships between words grounded in patterns of use, rather than linguistic theory. The high dimensionality of vector spaces in LLMs makes it possible to represent complex, multifaceted similarities between words—some of which plausibly reflect conceptual structure and/or inferential roles (Piantadosi2022).
@@ 4.4 Communicative Grounding
The notion of grounding is also used in relation to another aspect of language, having to do with communication (clarkGroundingCommunication1991, traumComputationalTheoryGrounding1994). In this context, the ‘grounding problem’ refers to the problem of establishing common ground between speakers in conversation. From this perspective, linguistic communication can be seen as a form of coordinated action that involves collaborating to reach a common understanding of what is said. The process of communicative grounding may involve explicit clarification strategies, such as prompting an interlocutor to repeat or reformulate a garbled or ambiguous statement. It also involves tracking communicative intentions during the exchange, to infer what is communicated from what is said. Such inferences can be made partly on the basis of pragmatic presuppositions, namely, what speakers take for granted or assume to be true when they use certain sentences in the context of the conversation (stalnakerCommonGround2002).
The communicative notion of grounding has also been discussed by natural language processing (NLP) researchers (brennan1998grounding, dimaroComputationalGroundingOverview2021, chanduGroundingGroundingNLP2021), and is closely tied to the challenge to LLM meaning advanced by Bender2020, who define meaning as a relation between linguistic expressions and communicative intent, characterising this relation as a form of grounding.
While communicative and referential grounding are distinct notions, they are related in important ways. Referential grounding anchors language to the world, but the same linguistic expressions may be referentially grounded in different ways for different speakers, due to differences in perception, knowledge, and conceptualisation. This means that calibrating reference in conversation requires communicative grounding—language users must work together to negotiate a common way of connecting language to the world. Without such collaboration, there is no guarantee that speakers really mean the same thing, even if they are using the same words.
@@ 4.5 Epistemic Grounding
Finally, the notion of grounding is occasionally used by NLP researchers to refer more minimally to the relationship between linguistic expressions and information stored in knowledge bases (poonGroundedUnsupervisedSemantic2013, tsaiConceptGroundingMultiple2016, chanduGroundingGroundingNLP2021, thoppilanLaMDALanguageModels2022). For example, an expert system may contain a knowledge base of facts and rules belonging to a particular domain. When the system receives user input, it can query this knowledge base to determine an appropriate response. A classic example is MYCIN, an expert system for diagnosing infectious blood diseases (shortliffeMycinKnowledgeBasedComputer1977). When given patient symptoms, MYCIN would search its knowledge base of medical rules and facts to suggest possible diagnoses and recommend treatments. In this system, language is grounded in the structured knowledge contained in the expert system’s knowledge base. We call this notion epistemic grounding.
There are two ways in which current LLMs may involve epistemic grounding. Firstly, they can be hooked up to external databases used to retrieve information (borgeaudImprovingLanguageModels2022, thoppilanLaMDALanguageModels2022). Secondly, it has been suggested that even LLMs that are not augmented with retrieval capabilities might be functionally similar to knowledge bases themselves (petroniLanguageModelsKnowledge2019). Indeed, they may encode a significant amount of world knowledge in their weights after pre-training, as illustrated by the remarkable performance of state-of-the-art LLMs on a wide range of factual question-answering benchmarks (openaiGPT4TechnicalReport2023, bubeckSparksArtificialGeneral2023).
@@ 4.6 What Grounding for LLMs?
We have identified five distinct notions that play prominent roles in discussions about grounding in biological and artificial systems. Which of these is necessary for LLMs to possess representations and produce outputs with intrinsic meaning? (Recall that intrinsic meaning is meaning that is independent from external interpretation.) To identify the most suitable notion, we will proceed by elimination, examining each candidate’s adequacy—beginning with the two kinds of sensorimotor grounding.
The ‘sensorimotor contact’ version of sensorimotor grounding claims that causal contact between sensorimotor transducers and the world grounds downstream representations. While promising, this view does not explain how the brute impinging of light rays, sound waves and the like on sensory surfaces can ground representations in the specific worldly things they are about. Indeed, such representations are typically not about light rays or sound waves, but rather about horses, democracies, elms, and so forth. [^14] Mutatis mutandis for electric signals that activate muscles vis-à-vis motor plans, actions, and behaviours. By itself, this sort of causal contact with the world does not explain how internal representations come to be about the latter, rather than (just) the former. As we will see (section 5), an account of how sensorimotor transducers can contribute to grounding requires additional components to be in place, making the sensorimotor contact version of sensorimotor grounding collapse into the notion of referential grounding: the causal contact with the world that sensorimotor transducers allow is just one of the components that can underlie referential grounding. As we will argue in section 6, moreover, such unmediated causal contact with the world is not even necessary for referential grounding.
The other version of sensorimotor grounding, the ‘sensorimotor representation’ view, connects conceptual or linguistic representations to sensorimotor representations within the same cognitive system. While this connection can illuminate the relevant representations’ role within a cognitive architecture, it fails to address the merry-go-round problem. By grounding one type of representation in another, it merely postpones the crucial question: how do sensorimotor representations themselves establish contact with the world? Since sensorimotor representational grounding does not provide a connection with the world but only with other representations, it does not address the Symbol Grounding Problem. Similarly, it offers no resolution to the Vector Grounding Problem, as grounding vectors in other vectors leaves unanswered how any vector ultimately connects to worldly entities and properties.
@@
This line of reasoning extends to relational, communicative, and epistemic grounding. Relational grounding anchors linguistic representations to their relationships to other linguistic representations. Although this kind of grounding clearly does not escape the representational carousel, it is plausible that it can suffice for endowing at least some representations with intrinsic meaning. Consider, for instance, logical connectives, for which an inferential-role semantics is compelling (blockAdvertisementSemanticsPsychology1986, Fodor1990). The meaning of a term like ‘and’ is arguably exhausted by the role it plays in licensed inferences; it does not purport to refer to a worldly entity, but to an abstract relation between propositions. Insofar as an LLM can induce the correct inferential role for logical terms from its training data, its internal states corresponding to these terms could be said to be meaningfully grounded in that role. [^15] Therefore, if an LLM can learn the structure of relations that constitute the meaning of logical connectives and linguistic conjunctions from its training data, then its representations of these terms might be as intrinsically meaningful as they can be. However, logical connectives are a rather special case; at best, inferential-role semantics accounts for only a narrow subset of the representations and outputs that biological and artificial systems possess and produce.
A similar case can be made for the representation of purely syntactic properties. The function of an internal state that represents, say, a subject-verb agreement rule or a filler-gap dependency is exhausted by its contribution to generating grammatically well-formed token sequences. Its correctness conditions are defined entirely within the formal system of the language itself. There is strong empirical evidence that LLMs, through the objective of next-token prediction alone, do in fact develop internal mechanisms that implement such grammatical rules. [^16] See hardingOperationalisingRepresentationNatural2023c for a discussion of criteria for ascribing representation of syntactic features to LLMs, and Milliereforthcoming for empirical evidence that LLMs do learn such representations. However, the Vector Grounding Problem is not about whether a system can acquire representations of intra-linguistic structures—be they logical or syntactic—but whether it can acquire representations with extra-linguistic content. Conceding that LLMs can achieve relational grounding for formal aspects of language does nothing to solve the problem of how their representations could come to be about worldly entities and properties. Thus, while relational grounding may well be sufficient for a narrow and special subset of representations whose content is exhausted by their formal role, it remains insufficient for the vast lexicon of terms that purport to refer to the world. For these terms, relational grounding merely spins the wheels on the representational merry-go-round; it cannot, by itself, provide the traction needed to connect with extra-linguistic reality.
@@
In sum, the form of grounding that underlies the others and directly addresses the Grounding Problems is what we have called referential grounding. Unlike the alternatives, referential grounding anchors representations in the world, rather than in other representations. It alone breaks free from the representational merry-go-round by connecting representations to whatever worldly entity or property they are about. Therefore, to determine whether generative pre-trained models like LLMs can manipulate and produce representations with intrinsic meaning, we must assess their capacity for referential grounding. This assessment requires examining what conditions must be met for a representation to achieve referential grounding.
@@ 5 Theories of Representational Content: Causes and History
Some internal representations are basic, in the sense that their meaning, or content, does not derive from other states already endowed with meaning. They can be seen as the building blocks of more complex representations, which draw some of their meaning from those basic representations. For instance, ‘edge detectors’ in primary visual cortex (V1) are good candidates for being basic internal representations, whereas ‘shape detectors’ in downstream visual areas are not, since their contents derive partly from the contents of the basic edge detectors. Similarly, words and sentences are not basic in this sense, since their meaning derives from other, more basic representational states, such as beliefs, desires and intentions—themselves complex representations.
The distinction between basic and complex representations cuts across the distinction between intrinsic and extrinsic meaning. Both the representations in primary visual cortex and the words and sentences humans utter are cases of representation with intrinsic meaning, though the former are basic while the latter are complex. By contrast, words randomly formed by the movement of ants in the sand or outputs from a random sentence generator only have extrinsic meaning: they are merely shapes that require human interpretation to convey any meaning at all.
We cannot here do justice to the rich and nuanced interdisciplinary research on the notion of cognitive representation. [^17] See Neander2017, Millikan2017, Shea2018 for recent book-length defences of the main contenders today. What matters for our limited purposes is that, despite some disagreements on the finer details, philosophers of cognitive science have largely reached a consensus on the key factors that enable internal states in a system to become representations, both basic and complex (Shea2018, CoelhoMollo2022).
@@
Two natural relations are widely considered to be central to endowing basic internal states with meaning: causal-informational relations, and historical relations. In general terms, causal-informational relations obtain when one state of the world conveys information about another due to a causally-generated correlation between the two states. A classic example is the relationship between smoke and fire: smoke is correlated with fire because it is caused by it, so the presence of smoke carries information about the presence of fire. Neurons in brain area V1 convey information about edges in the perceived visual scene, as their activation correlates (through a relatively complex causal chain that includes light rays impinging on the retina) with the presence of edges in the world in front of the perceiving subject.
@@
Causal-informational relations capture the main rationale for why systems possess basic internal representations: they allow systems to be sensitive to specific features of their environments. However, the mere existence of a causal-informational relation is insufficient to provide an account of basic representational content. An additional requirement often mentioned is that the causal-informational relation be exploited, that is, used by the system due to its information-carrying nature (Shea2018). In other words, a system must use the presence of smoke to obtain information about the presence of fire, form a belief about there being a fire, guide fire-related behaviour, and so on, for smoke to be a viable candidate for representing fire for that system.
@@
To be a representation necessarily involves playing a certain role: that of conveying information about the world and being used by the system in virtue of that. This further requirement complements the explanation for why internal representations are useful to systems: they are used by the system as sources of information about the world. It also helps to avoid the risk of trivialising the notion. Causal-informational relations are abundant in nature: every effect carries information about its cause. If every such relation should suffice for being a representation, representations would be so trivial that their special explanatory power for cognition would be lost.
@@
Exploited causal-informational relations alone, however, are still not enough to base an adequate theory of representation, for another crucial feature of internal representations is missing: their capacity for truth or falsehood, correctness or incorrectness. [^18] Correctness conditions are typical of so-called descriptive representations. Directive representations, e.g., desires and action plans, are in the business of representing how aspects of the world should be. Such representations have satisfaction conditions, i.e., the conditions that would satisfy, say, a specific desire. For simplicity’s sake, we will be concerned only with descriptive representations. To be a representation inherently involves the possibility of misrepresenting. V1 neurons may become active when no edge is present, as in the famous Kanizsa’s triangle illusion, thus conveying information to the system that, if not corrected by other subsystems, can lead to inappropriate behaviour (such as trying to touch the non-existent edge). Representations involve a kind of descriptive normativity (Neander2017). That is to say, representations can represent successfully or unsuccessfully, better or worse. This descriptive normativity extends to external representations, such as sentences, and meaning more generally. They are kinds of things that can be true or false, accurate or inaccurate, appropriate or inappropriate. The rationale for systems to possess representations is that they be accurate enough, enough of the time. Otherwise, representations would lead systems astray and hinder appropriate behaviour, rather than the other way around. Generally, (descriptive) representations have the function to represent correctly and accurately (to a degree of required precision). [^19] In some special cases, systematic misrepresentation can be a sensible price to pay. If the stakes are high enough, it makes sense for representations to err on the side of caution: if you are trying to detect predators, it makes sense to produce many false-positive (misrepresenting predators as present when there are none), rather than even one, potentially fatal false-negative.
@@
As the notion of representation is central for cognitive (and computer) science, the notion of function is central to biology. This has led philosophers of biology to develop scientifically-grounded, objective theories of function (Wright1973, Millik1989, Neander1991, Godfrey-Smith1994, Garson2019). If functions exist out there in the world, they must be a matter of natural features and relations, rather than human interpretation. The dominant view of functions is the selected-effects theory (Millik1989, Neander1991, Garson2019). On this view, functions are ensembles of the historical causes that explain the persistence of a certain trait or feature in a system. For instance, hearts today have the function to pump blood because past hearts, by pumping blood, were selected due to their contribution to the survival and reproduction of the systems having them. Pumping blood is what explains, through a historical selection process, what hearts were selected for doing—and it is thereby their function. The relevant selection processes can be at the level of the species, typically through natural selection; or at the level of the individual, through selection-like learning processes during development.
@@
The most influential views on representation incorporate the selected-effects theory to account for internal states having the function to carry information about the world and to guide appropriate behaviour (Neander2017, Millikan2017, Shea2018, CoelhoMollo2022). The basic idea is that some states and processes in cognitive systems were selected for in virtue of their carrying information about, and guiding successful behaviour toward, states of the world—thus acquiring functions to represent such states. For example, neurons in V1 have the function of detecting edges in the visual scene because they were selected by natural selection in virtue of the causal-informational relations they maintain with actual edges in the world. When these neurons successfully detect edges, they fulfil their function, representing correctly; when they fail to fire in the presence of an edge or fire in the absence of one, they malfunction, misrepresenting. In brief, internal representations are states that carry information about the world, have the selected function to do so, and are appropriately used by the system in virtue of the information they carry.